# Singular basins in multiscale systems

### v1

@@ Abstract
Real-world complex systems often evolve on different timescales and possess multiple coexisting stable states. Whether or not a system returns to a given stable state after being perturbed away from it depends on the shape and extent of its basin of attraction. In this Letter, we show that basins of attraction in multiscale systems can exhibit special geometric properties in the form of universal singular funnels. We use the term singular basins to refer to basins of attraction with singular funnels. We show that singular basins occur robustly in a range of dynamical systems: the normal form of a pitchfork bifurcation with a slowly changing parameter, an adaptive active rotator, and an adaptive network of phase rotators. Although singular funnels are narrow, they can extend to different parts of the phase space and, unexpectedly, impact the resilience of the system to disturbances. Crucially, the presence of a singular funnel may prevent the usual dimensionality reductions in the limit of large timescale separation, such as quasi-static approximation, adiabatic elimination or time-averaging of the fast variables
@@
Complex real-world systems are characterized by multistability, i.e. when given different initial conditions, the same system can end up with notably different asymptotic behavior. Examples of multistability can be found in epileptic and neuronal models [jirsa_personalised_2023, Gerstner2014, Pfeifer2021, Minosa], reservoir computers [Flynn2021a], lasers [Jaurigue2017, Wieczorek2005, Yanchuk2019], and climate [alkhayuon2019basin, quinn2018mid, slyman2025tipping], to name a few. The analysis and control of multistability-related phenomena has therefore been the subject of many studies [Pisarchik2014, Ebrahimzadeh2022, Malashchenko2011, Hellmann2018]. An important question in multistable systems concerns the interplay between different stable states and possible transitions between them. In the context of cell differentiation, for instance, transitions between attractors by a perturbation correspond to the reprogramming of cells [huang_reprogramming_2009]. One notable approach addressing the above challenges is the concept of basin stability [Menck2013, soliman_global_1992], which represents the likelihood of reaching a particular stable state starting from random initial conditions. This likelihood can be estimated by the volume of the basin of attraction of a stable state, in other words, the proportion of initial conditions in the phase space that evolve into that state over time. More importantly, the geometry of the basins of attraction determines the resilience of a system to perturbations [holling_resilience_1973, krakovska_resilience_2024]. Therefore, studying basins of attraction of different stable states, whether their relative volumes or, more broadly, their geometries, is crucial for understanding and controlling multistable systems.
@@
In addition to having multiple stable states, it is also common for real-world systems to evolve on multiple timescales, which give rise to a variety of nonlinear phenomena [Kuehn2015, Krupa1997, balzer_canard_2024, ritchieRateinducedTippingNatural2023, Desroches2012]. An important example of a multi-scale system is an adaptive dynamical network, where the adaptation is much slower than the node dynamics [Berner2023, Berner2021, Fialkowski2023, Maslennikov2018, gross2008adaptive, Aoki2011]. Such systems are challenging because they are high-dimensional, and can exhibit high multistability with a large number of coexisting stable states [Rohr2019]. The study of the basins of attraction for multiscale systems is a largely unexplored area of research, and revealing the properties of the basins arising from the interplay of multiple timescales is an intriguing and important task.
A number of useful techniques have been developed for analysing mathematical models of multiscale systems with two distinct timescales, also known as slow-fast systems [jones_geometric_1995, wechselberger_geometric_2020, Fenichel1979, krupaRelaxationOscillationCanard2001]. These include adiabatic elimination, quasistatic approximation, and averaging, which allow the elimination of the fast variables by taking into account their effective averaged contribution to the slow dynamics [hummel_reduction_2023, hasselmann_stochastic_1976, Politi2023].
@@
This Letter reports specific properties of basins in slow–fast systems that arise from their multiscale separation. We show that these basins can contain singular funnels (SFs) in the form of tunnels that become increasingly narrow as the timescale separation grows. Due to this vanishing property, we refer to basins containing SFs as singular basins. In systems with SFs, adiabatic elimination or the quasistatic approximation causes the SFs to disappear and, consequently, eliminates the possibility of reaching a given steady state from certain regions of phase space. Therefore, one must be cautious when extrapolating the resilience properties of the full system from those of a reduced one, whether obtained via quasistatic approximation, adiabatic elimination, or averaging.
For example, if an averaged system has a bistable potential with two stable states, the basins of attraction of these states are clearly separated. In other words, there exists a threshold value of the slow variable such that, on each side of it, the system evolves toward a different steady state. However, when the fast subsystem is taken into account, SFs enable convergence to the same steady state from both sides of the threshold value of the slow variable, even when the separation of timescales is arbitrarily large. From the perspective of the reduced system, this appears as a “tunnelling” from one state to another, Fig. 1(b).
We show the singular basin in different classes of systems: a pitchfork bifurcation normal form with a slowly changing parameter, an active rotator with an adaptive frequency, and an adaptive dynamical network of active rotators. We further show that the phenomenon is robust to parameter changes and exhibits universal scaling with increasing timescale separation in small dimensions. We conclude that singular basins are common in multiscale systems, and specify the main conditions under which this phenomenon occurs.
We first present the singular basin in a normal form for the pitchfork bifurcation, where the bifurcation parameter is slowly changing. The pitchfork bifurcation is one of the few generic mechanisms for stability exchange in dynamical systems, and its normal form equation is $\dot{x}(t)=x(\mu-x^{2})$ in the supercritical case [Kuznetsov1995]. Here $x$ is a dynamical variable and $\mu$ a parameter. We restrict the normal form system to non-negative $x$ values, which is appropriate for many applications, for example, when the variable $x$ denotes a population size or laser intensity. Then, the system has a single stable steady state for all parameters $\mu$, $x^{*}=0$ for $\mu\leq 0$, and $x^{*}=\sqrt{\mu}$ for $\mu\geq 0$. In a more compact form, one can write the stable equilibrium as $x^{*}=\sqrt{H(\mu)\mu}$, where $H(\mu)$ is the Heaviside step function $H(\mu)=0$ for $\mu<0$ and $H(\mu)=1$ for $\mu\geq 0$. The slow dynamics is given by a linear adaptation of the bifurcation parameter $\mu$, resulting in the following system
@@
where $\varepsilon$ is a small positive parameter, and $a,b>0$ are parameters determining the adaptation rule. Under the condition $a>2\sqrt{b}$, system (12) possesses three equilibria: two stable equilibria $e_{0}$ and $e_{2}$, and one saddle point $e_{1}$ as shown in Fig. 1(a), see calculations in Supplemental Material [supplem].
Figure 1: (a-b): Singular basin and singular funnel (SF) in a pitchfork normal form (12) with slowly adapting parameter $\mu$. (a) Phase space of the system with two coexisting stable equilibria $e_{0}$ and $e_{2}$ and the corresponding basins of attraction. The red (white) region is the basin of attraction of $e_{0}$ ($e_{2}$). SF is observed near the $x=0$ axis, it extends to arbitrarily large $\mu$ values and decreases exponentially in width. Parameters: $a=3$ , $b=2$, and $\varepsilon=0.1$. (b) The bistable potential of the corresponding slow system after adiabatic elimination of the fast variable (equivalently, reduction to the stable critical manifold). (c-d): Same for system (26). Parameters: $a=5$ , $b=10$, and $\varepsilon=0.1$.
@@
Since the fast subsystem (1) possesses a unique stable equilibrium $x^{*}(\mu)=\sqrt{H(\mu)\mu}$ for all values of the slow variable $\mu$, we apply the adiabatic elimination procedure, i.e., substitute $x^{*}(\mu)$ into (2) and obtain the reduced slow system
$\dot{\mu}=\varepsilon f(\mu),$ (3)
with $f(\mu)=-\mu+a\sqrt{H(\mu)\mu}-b$ and the corresponding potential $U(\mu)=-\int f(\mu)d\mu=\mu^{2}/2+b\mu-2a\mu\sqrt{H(\mu)\mu}/3$ as shown in Fig. 1(b). Hence, the adiabatic elimination reduces system (12) to a simple bistable potential system for the slow variable $\mu$.
In the reduced system (3), the fast dynamics is eliminated and the attraction basins of the stable equilibria $\mu_{0}$ and $\mu_{1}$ are separated by $\mu_{2}$. Therefore, all states of (3) initially started with $\mu<\mu_{2}$ tend to $\mu_{0}$ over time, while all $\mu>\mu_{2}$ tend to $\mu_{1}$. This clear division of basins by $\mu=\mu_{2}$ is not present in the original slow-fast system (12) even for an arbitrary large time-scale separation $1/\varepsilon$. Figure 1(a) shows the basins of attraction of the stable states $e_{0}$ and $e_{2}$. It can be seen that states with arbitrarily large $\mu$ can converge to the left equilibrium $e_{0}$ over time when the initial conditions are chosen in the SF, which is a small stripe close to $x=0$. This funnel-like part of the attraction basin has an exponentially small width as $\varepsilon$ decreases, leaving the possibility of "tunneling" towards the $e_{0}$ state. The SF clearly restricts the applicability of the adiabatic elimination.
@@
One might suspect that the failure of a global adiabatic elimination is caused by the degenerate point at $(\mu,x)=(0,0)$ in the pitchfork normal form (12). However, to demonstrate that this is not the case, we present the singular basin for a modified model:
@@
$\dot{\mu}(t)$ $=\varepsilon(-\mu+ax-b)$
in Fig. 1(c-d). The singular basin appears there without any degenerate point.
An exponentially narrow SF observed in the paradigmatic models (12) and (26) appears to be characteristic for more complex classes of slow-fast systems. We will demonstrate this for an adaptive phase rotator model and an adaptive network of coupled phase rotators.
The fast-slow adaptive phase rotator has the form [Franovic2020]
@@
$\dot{\mu}(t)$ $=\epsilon(-\mu+\eta(1-\sin(\varphi+\alpha))),$ (6)
where $\varphi$ is a phase variable and $\mu$ accounts for a slow self-adjustment of the oscillator frequency. Although system (56) has the same dimension as the normal form (12), the fast variable $\varphi$ can describe periodic rotations. In particular, the fast subsystem (5) has a pair of stable and unstable equilibria for $|\omega+\mu|<1$ and a rotating solution for $|\omega+\mu|>1$ [^1] This corresponds to the case of unboundedly growing $\varphi(t)$ when $\varphi$ considered on a real line $\mathbb{R}$ or, equivalently, rotating on a circle $\mathbb{S}^{1}$ when $\varphi$ considered modulo $2\pi$ phase. At the point $|\omega+\mu|=1$, the fast subsystem undergoes a "saddle-node on an invariant curve" bifurcation [Franovic2020, Strogatz1994]. Hence the fast subsystem (5) has a single stable attractor for all fixed values of $\mu$, which is an equilibrium for $|\omega+\mu|\leq 1$ and a rotation for $|\omega+\mu|>1$.
The adaptive phase oscillator (56) can also be reduced to the form (3) after an elimination of the fast timescale. This is achieved by averaging over the fast rotations for $|\omega+\mu|>1$ and by adiabatic elimination for the remaining values of $\mu$ on the single stable equilibrium. The corresponding function $f(\mu)$ of the reduced system (3) can be calculated explicitly:
@@
The details of the derivation can be found in the Supplemental Material [supplem].
@@
We consider the case when system (56) possesses two coexisting attractors: a stable equilibrium $e_{1}$ and a stable rotation $\gamma_{c}$, see Fig. 2(a,b). This bistability occurs for a large set of parameters $(\varepsilon,\alpha,\eta)$, which is described in detail in Supplemental Material [supplem]. The reduced system (3), with the right-hand side given by equation (7), is also bistable in this case, and its potential is shown in Fig 2(d). Here $\mu_{1},\mu_{c},$ and $\mu_{2}$ are the $\mu$-coordinates of the equilibrium $e_{1}$, the rotation $\gamma_{c}$ in the limit $\varepsilon\to 0$, and the saddle point $e_{2}$, respectively.
@@
Figure 2: Singular basin and singular funnels in an adaptive phase-oscillator (56). (a) and (c): The red shaded area is the basin of attraction of the limit cycle $\gamma_{c}$, the white area is the basin of attraction of the equilibrium $e_{1}$. The boundary of the two basins is given by the stable manifold of the saddle equilibrium $e_{2}$ (blue lines). (b): The bistable potential of the reduced system (3,7) after averaging and adiabatic elimination of the fast variable. (d) The volume of the singular funnel of $\gamma_{c}$ for $10\leq\mu\leq 0$ as a function of $\varepsilon$. Red circles: the volume computed using Monte Carlo simulation [supplem]. Black line: the volume as given by equation (8). Other parameters: $\eta=10$, $\omega=-4$. (a): $\alpha=\pi/2$, $\varepsilon=0.1$, (c): $\alpha=\pi/2$, $\varepsilon=0.01$.
@@
The singular basin of the adaptive phase oscillator (56) is shown in Figures 2(a) and (b). The SF has an exponentially decreasing width for small $\varepsilon$ and extends into the region of negative $\mu$. This creates a channel that allows for initial conditions with arbitrary $\mu$ and selected $\varphi$ to be attracted to the rotating attractor $\gamma_{c}$, which is impossible in the framework of the averaged system (3).
@@
Next, we provide details of how the boundaries of the SF are formed in Fig. 2(a,b). As in the pitchfork normal form case (12), these boundaries consist of orbits attracted to the saddle equilibrium $e_{2}$ (the stable manifolds). Tracking these orbits backwards in time reveals that they pass close to the unstable equilibria of the fast system (the dashed part of $S$ in Fig. 2(a,b)), which is also known as the critical manifold [Kuehn2015]. Since this part of the critical manifold is stable in backward time, the two SF boundaries are exponentially attracted, causing the SF to become exponentially narrow. After the SF has passed $S$, it undergoes further rotations and extends to arbitrarily negative $\mu$, maintaining the same exponentially narrow width.
@@
For the considered systems with singular basins, we can obtain the scaling of the SF phase volume as a function of $\varepsilon$ and estimate how quickly it vanishes as $\varepsilon$ approaches 0. In all cases, the motion along the trajectories, which correspond to the SF boundaries, is slow and its speed is proportional to $\varepsilon$ along the critical manifold $S$, so the time they spend close to a finite piece of length $L$ of $S$ scales as $L/\varepsilon$. The width $\delta$ of the SF is determined by two trajectories that pass close to $S$ (blue trajectories in Fig. 2) and, hence, $\delta$ decreases exponentially as time goes backwards. Therefore, this width can be estimated as $\delta\sim\exp(-L\lambda/\varepsilon)$ at the moment when the boundaries leave the critical manifold $S$ (or a section of it of length $L$). Here $\lambda$ is an effective repulsion rate at $S$, which corresponds to the attraction in the reverse time. These arguments lead to the following expected scaling for the volume of a SF:
@@
The scaling (8) clearly holds for the normal form system (12) and system (26). Furthermore, the rotator system (56) exhibits this scaling too, as shown in Fig. 2(c). The basin scaling (8) is expected to hold for large classes of slow-fast systems, at least in low dimensions, as the arguments leading to (8) are based on the rather general geometric structure of the basin boundaries. However, as we will see below, this scaling may sometimes fail to accurately describe singular basins in high-dimensional systems.
Singular basin is a robust phenomenon. This is due to the fact that its ingredients persist under small parameter changes. This applies to the stable manifolds of the saddle equilibria and the coexisting attractors. We illustrate the parameter region of the singular basin and the possible mechanisms of its destruction for system (56) in the Supplemental Material [supplem].
@@
where $i=1,\dots,N$ and $X=\frac{1}{N}\sum_{j=1}^{N}\sin\left(\varphi_{j}+\alpha\right)$. In this system, the frequencies of the individual rotators are adapted globally by the slow variable $\mu$, and $\mu$ is driven by the mean-field $X$. This could be viewed as an interaction with the environment. In contrast to the extensively studied system of coupled rotators without adaptation [Park1996, Lindner2004, Dolmatova2017, Klinshov2021, Ronge2021, burylkoTimereversibleDynamicsSystem2023, Hellmann2018, Schafer2018], the adaptive system (910) contains multiple timescales. Adaptive systems of a similar nature have been shown to exhibit distinct dynamical properties such as canard cascading [DHuys2021, balzer_canard_2024], emergent excitability [Ciszak2020a], and others [Song2020, ciszak_collective_2021, venegas2023stable].
Figures 3 and 4 show singular basins and their scaling in systems of $N=2$ and $N=10$ rotators, respectively. Both systems feature the coexistence of two attractors: one stable equilibrium and another rotating periodic or quasi-periodic attractor. Consequently, the corresponding averaged system exhibits a bistable potential, as shown in Fig.3(a). This bistability is caused by the phase-space structures similar to those in the single adaptive rotator. However, a detailed description of these structures is beyond the scope of this letter.

### v2

@@ Abstract
Real-world systems often evolve on different timescales and possess multiple coexisting stable states. Whether or not a system returns to a given stable state after being perturbed away from it depends on the shape and extent of its basin of attraction. We show that basins of attraction in multiscale systems can exhibit special geometric properties in the form of singular funnels. Although singular funnels are narrow, they can extend to different regions of the phase space and, unexpectedly, impact the system’s resilience to perturbations. Consequently, singular funnels may prevent common dimensionality reductions in the limit of large timescale separation, such as the quasi-static approximation, adiabatic elimination and time-averaging of the fast variables. We refer to basins of attraction with singular funnels as singular basins. We show that singular basins are universal and occur robustly in a range of multiscale systems: the normal form of a pitchfork bifurcation with a slowly adapting parameter, an adaptive active rotator, and an adaptive network of phase rotators.
@@
Complex real-world systems are characterized by multistability, i.e. when started from different initial conditions, the same system can end up with notably different asymptotic behavior. Examples of multistability can be found in epileptic and neuronal models [27, 19, 42, 38], reservoir computers [17], lasers [26, 56, 57], and climate [1, 45, 49], to name just a few. The analysis and control of multistability-related phenomena has therefore been the subject of many studies [43, 14, 37, 22]. An important question in multistable systems concerns the interplay between different stable states and possible transitions between them. For example, in the context of cell differentiation, such transitions correspond to the reprogramming of cells [24]. One notable approach addressing the above challenges is the concept of basin stability [40, 50], which represents the likelihood of reaching a given state (attractor) when starting from a random initial condition. This likelihood can be estimated by the relative volume of its basin of attraction, i.e., the proportion of initial states in the phase space that converge to this attractor. More importantly, the shape of the basins of attraction determines the uncertainty [11] and resilience [23, 30] of a system to perturbations. Therefore, to understand and control a multistable system, it is crucial to understand the geometry of its basins of attraction.
@@
In addition to having multiple stable states, it is also common for real-world systems to evolve on multiple timescales, which gives rise to a variety of nonlinear phenomena [34, 33, 4, 46, 12]. An important example of a multi-scale system is an adaptive dynamical network, where the adaptation is much slower than the node dynamics [5, 6, 16, 39, 20, 2, 54]. Such systems are challenging because they are also high-dimensional and can exhibit a high degree of multistability. The study of basins of attraction in multiscale systems is an area of research that remains largely unexplored. Therefore, uncovering new properties of the basins that arise from the interplay of different timescales is an intriguing and important task.
Useful and widely used techniques have been developed to analyze mathematical models of multiscale systems with two distinct timescales, also known as slow-fast systems [28, 55, 15, 32]. These techniques exploit the limit of infinite ratio between the fast and slow timescales and can be broadly divided into two groups [25, 21, 44, 34]. The first group includes adiabatic elimination and averaging. It eliminates the fast variables by treating them as instantaneous, or by taking into account their effective contribution, while only considering the time evolution of the slow variables in the reduced slow subsystem. The second group includes the quasistatic approximation. It ‘freezes’ the slow variables by treating them as static at each instant, while only considering the time evolution of the fast variables in the reduced fast subsystem.
@@
This Letter uncovers special properties of basins of attraction in slow–fast systems, leading to counter-intuitive phenomena. We show that such basins can contain singular funnels (SFs) in the form of tunnels that extend to different regions of the phase space and become increasingly narrow as the timescale ratio increases. The problem with SFs is that they are eliminated by the quasistatic approximation and adiabatic elimination. Consequently, adiabatic elimination makes it impossible to reach a given stable state from certain regions of the phase space, whereas this is possible in the full system. On the other hand, the quasistatic approximation could mask the presence of multiple basins of attraction entirely. Therefore, one must be cautious when extrapolating the resilience properties of the full system from those of a reduced one, regardless of whether the latter is obtained via quasistatic approximation, adiabatic elimination, or averaging. Due to the vanishing of SFs in the limit of infinite ratio of the timescales, we refer to basins containing SFs as singular basins.
[…]
@@
where the dot denotes $d/dt$. The small parameter $0<\varepsilon\ll 1$ is the ratio between the fast timescale of $x$ and the slow timescale of $\mu$. When $\varepsilon=0$, the slow timescale is infinitely slower, or equivalently, the fast timescale is infinitely faster. The parameters $a,b>0$ determine the adaptation rule. Under the condition $a>2\sqrt{b}$, system (12) has two stable equilibria, $e_{1}$ and $e_{3}$, one saddle equilibrium, $e_{2}$, and a SF, as shown in Fig. 1(a); see the supplemental material for more details.
@@
The fast subsystem is obtained by setting $\varepsilon=0$ in system (12), which gives system (1) with static $\mu$. Since it has a unique stable quasistatic equilibrium $x^{*}(\mu)$ that varies continuously with $\mu$ (the solid part of the gray curve $S$ in Fig. 1(a)), we apply the adiabatic elimination procedure. We substitute $x^{*}(\mu)$ into (2) and obtain the reduced slow subsystem
$\frac{d\mu}{d\tau}=f(\mu),$ (3)
where $f(\mu)=-\mu+a\sqrt{\mu H(\mu)}-b$ and $\tau=\varepsilon t$ the slow time. We then recognise that this is a bistable dynamical system with a double-well potential $U(\mu)=-\int f(\mu)d\mu=\mu^{2}/2+b\mu-2a\left(\mu H(\mu)\right)^{3/2}/3$ and the same equilibria as the full system; see Fig. 1(b). After adiabatic elimination, the basins of attraction of the stable equilibria $e_{1}$ and $e_{3}$ in Fig. 1(b) are clearly separated by a single value of the slow variable, $\mu=\mu_{b}$, given by the position of the unstable equilibrium $e_{2}$. If the slow subsystem is started on the left (resp. right) of $\mu_{b}$, it can only converge to $e_{1}$ (resp. $e_{3}$). However, this clear separation of basins by $\mu=\mu_{b}$ is never present in the full system (12) due to the SF, no matter how large the timescale ratio is (or how small $\varepsilon$ is). In particular, the SF allows convergence to $e_{1}$ from both sides of $\mu_{b}$, which is impossible in the reduced slow subsystem. This includes perturbations from $e_{3}$ into the SF along the fast $x$-direction. From the slow subsystem’s perspective, this appears as though there is quantum-like ‘tunnelling’ or ‘teleporting’ from $e_{3}$ to $e_{1}$ (the grey arrow in Fig. 1(b)). In summary, the SF clearly restricts the applicability of the adiabatic elimination, even when the fast subsystem has a unique stable quasistatic equilibrium that varies continuously with $\mu$.
Figure 1: (a) Singular basin with a singular funnel (SF) in a bistable pitchfork normal form (12) with slowly adapting parameter $\mu$. The red (white) region is the basin of attraction of the stable equilibrium $e_{1}$ ($e_{3}$). The boundary of the two basins is given by the stable manifold of the saddle equilibrium $e_{2}$ (blue curves). The SF near $x=0$ shrinks exponentially with $\mu$. (b) The bistable potential of the corresponding slow subsystem after adiabatic elimination of the fast variable $x$ (equivalently, reduction to the stable critical manifold). Parameters: $a=3$ , $b=2$, and $\varepsilon=0.1$. (c)-(d) The same as (a)-(b) but for system (26) with $a=5$ , $b=10$, and $\varepsilon=0.1$.
@@
One might suspect that the failure of a global adiabatic elimination for system (12) is caused by the degeneracy at $(\mu,x)=(0,0)$, where $x^{*}(\mu)$ is stable but non-hyperbolic. To demonstrate that this is not the case, we consider a modified system
@@
$\dot{\mu}(t)$ $=\varepsilon(-\mu+ax-b),$
that does not have any degeneracies. As shown in Fig. 1 (c–d), this system exhibits a singular basin with a SF and the same failure of global adiabatic elimination.
The SFs in the paradigmatic models (12) and (26) have very simple geometry. In general, however, SFs can have intricate geometries and can penetrate the (white) basin of the other stable state in complicated ways. We will now demonstrate this for adaptive phase rotators.
The fast-slow adaptive phase rotator has the form [18]
@@
$\dot{\mu}(t)$ $=\varepsilon(-\mu+\eta(1-\sin(\varphi+\alpha))),$ (6)
where $\varphi\in(0,2\pi]$ is a phase variable, $\omega$ is the base oscillator frequency, and $\mu$ accounts for a slow self-adjustment of the oscillator frequency. The fast subsystem (5) with static $\mu$ has a pair of stable and unstable equilibria for $|\omega+\mu|<1$, and a periodic rotation for $|\omega+\mu|>1$  [^2] This corresponds to the case of unboundedly growing $\varphi(t)$ when $\varphi$ considered on a real line $\mathbb{R}$ or, equivalently, rotating on a circle $\mathbb{S}^{1}$ when $\varphi$ considered modulo $2\pi$ phase. At the point $|\omega+\mu|=1$, the fast subsystem undergoes a "saddle-node bifurcation on a periodic rotation" [18, 52]. Hence, the fast subsystem (5) has a single stable attractor for all fixed values of $\mu$, which is an equilibrium for $|\omega+\mu|\leq 1$ and a periodic rotation for $|\omega+\mu|>1$.
The adaptive phase rotator (56) can also be reduced to the form (3) by eliminating the fast variable $\varphi$. This is achieved by averaging over the fast rotations for $|\omega+\mu|>1$ and by adiabatic elimination for $|\omega+\mu|<1$. The corresponding function $f(\mu)$ of the slow subsystem (3) can be calculated explicitly; see the supplemental material for more details:
@@
We consider the case when system (56) possesses two coexisting attractors: a stable equilibrium $e_{1}$ and a stable rotation $\gamma_{c}$; see Fig. 2(a) and the supplemental material for more details. The slow subsystem (3), with $f(\mu)$ given by equation (7), is also a bistable system with a double-well potential as shown in Fig 2(b).
@@
Figure 2: (a,c) Singular basin with a singular funnel (SF) in an adaptive phase-oscillator (56). The basin of attraction of the periodic rotation $\gamma_{c}$ is plotted in red and the basin of attraction of the stable equilibrium $e_{1}$ is plotted in white. The (blue) boundary of the two basins is given by the stable manifold of the saddle equilibrium $e_{2}$. (b) The bistable potential of the corresponding slow subsystem after adiabatic elimination and averaging of the fast variable $\varphi$. (d) The volume of the SF limited to $10\leq\mu\leq 3$ as a function of $\varepsilon$ obtained using equation (8) (black curve) and Monte Carlo simulation (red open circles) see the supplemental material for details. Other parameters: $\eta=10$, $\omega=-4$; (a): $\alpha=\pi/2$, $\varepsilon=0.1$; (c): $\alpha=\pi/2$, $\varepsilon=0.01$.
@@
The singular basin of the adaptive phase oscillator (56) in Fig. 2(a,c) has a SF that extends into the region of negative $\mu$. This creates a channel through which initial conditions with arbitrary $\mu$ and selected $\varphi$ are attracted to the periodic rotation $\gamma_{c}$. Once again, this is impossible within the framework of the averaged system (3), where only initial conditions with $\mu>\mu_{b}$ can converge to the averaged periodic rotation $\gamma_{c}$; see Fig. 2(b).
@@
Next, we provide details of how the boundaries of the SF in Fig. 2(a,c) are formed. As in the case of the adaptive pitchfork normal form (12), these boundaries consist of orbits attracted to the saddle equilibrium $e_{2}$ (branches of the stable manifold of $e_{2}$). Tracking these orbits backwards in time reveals that they pass close to the branch of unstable quasistatic equilibria of the fast subsystem (the dashed part of the curve $S$ in Fig. 2(a,c)), which is also known as the critical manifold [34]. Since this part of the critical manifold is attractive in backward time, the two SF boundaries exponentially approach each other as they are attracted to $S$, causing the SF to become exponentially narrow. After the SF extends past $S$ to negative $\mu$, it undergoes further rotations and repeatedly intersects the white basin of attraction of $e_{1}$ leading to a SF with a complicated pattern.
@@
For the systems with a singular basin that have been considered so far, we can obtain the universal scaling of the SF volume in phase space as a function of $\varepsilon$ to estimate how quickly SFs shrink as $\varepsilon$ approaches 0. In all cases, the motion along the trajectories, which correspond to the SF boundaries, is slow and its speed is proportional to $\varepsilon$ along the unstable branch of a critical manifold $S$. Thus, the time they spend close to an unstable branch of $S$ of length $L$ scales as $L/\varepsilon$. The width $\delta$ of the SF is determined by two trajectories that pass close to $S$ (blue trajectories in Fig. 2). Hence, $\delta$ decreases exponentially as time goes backwards. Therefore, this width can be estimated as $\delta\sim\exp(-L\lambda/\varepsilon)$ at the moment when the SF leaves the neighbourhood of an unstable branch of $S$ of length $L$. Here $\lambda$ is an effective repulsion rate at $S$, which corresponds to the attraction in the reverse time. These arguments lead to the following expected scaling for the volume of a SF:
@@
The scaling (8) is in perfect agreement with the adaptive normal form (12) and system (26). Furthermore, the adaptive phase rotator (56) exhibits this scaling too, as shown in Fig. 2(c). The SF scaling (8) is expected to hold for large classes of slow-fast systems, at least in low dimensions, as the arguments leading to (8) are based on the rather general geometric structure of the basin boundaries. However, as we show below, SFs with intricate geometries found in higher-dimensional systems may show deviations from this scaling.
Figure 3: (a) Bistable potential of the slow subsystem of two adaptive phase oscillators (9-10), for $N=2$, after adiabatic elimination and averaging of the fast variables $\varphi_{1,2}$. (b, d) Two cross-sections of the singular basin with a singular funnel (SF) of the same system: (b) $\varphi_{1}=1.2461$ and (d) $\mu=2.86$. The basin of attraction of the periodic rotation $\gamma_{c}$ is plotted in red and the basin of attraction of the stable equilibrium $e_{1}$ is plotted in white. The (blue) boundary of the two basins is given by the stable manifold of the saddle equilibrium $e_{2}$. (c) The volume of the SF limited to $10\leq\mu\leq 0$ as a function of $\varepsilon$ obtained using Monte Carlo simulation for different values of the parameter $\Delta_{\omega}=\omega_{1}-\omega_{2}$, showing the robustness of the SF. Other parameters: $\omega_{1}=-4$, $\kappa=1$, $\eta=10$; (b-d) $\varepsilon=0.1$.
@@
where $i=1,\dots,N$ and $X=\frac{1}{N}\sum_{j=1}^{N}\sin\left(\varphi_{j}+\alpha\right)$. In this system, the frequencies of the individual rotators are adapted globally by the slow variable $\mu$, and $\mu$ is driven by the mean-field $X$. In contrast to the extensively studied system of coupled rotators without adaptation [41, 36, 13, 29, 47, 7, 22, 48], the adaptive system (910) evolves on two distinct timescales with ratio $\varepsilon$. Adaptive systems of a similar nature have been shown to exhibit distinct dynamical properties such as canard cascading [10, 4], emergent excitability [8], and others [51, 9, 53]. Here, we add singular basins to this list of distinct properties.